Texas Products Inc. has a division that makes plastic composite bags for the space industry. The division has fixed costs of $45,000 per month, and it expects to sell 45,000 bags per month. If the variable cost per bag is $6.00, what price must the division charge in order to break even?
Answer by JP(22)

Ok first make a list of your given variables.
Cost:$45,000/month
Quantity sold:45,000 bags/month
Cost/bag:$6.00
First calculate how much it will cost to sell 45,000 bags/month at a purchase price of $6.00.
(45,000)(6.00)=$270,000 cost to sell the 45,000 bags at purchase price of $6.00/bag.
There is a fixed cost of $45,000 to I assume, run the company, so this cost must be included.
So $270,000+$45,000=$315,000 total expense
So you must now set up an equation...
x is equal to the amount of money you must sell your bags at in order to break even on expenses to profit ratio.
(45,000 bags)(x)=$315,000 expense
So, 45,000x=315,000
divide 315,000 by 45,000 to isolate the "x" variable and you get $7.00 as your solution.
x=315,000/45,000
x=7
To check your work, set up a function.
F(x)=45,000x=315,000
For every instance of x plug in your new solution
F(7)=(45,000)(7)=315,000
Then you get...
315,000=315,000 it proves the solution was correct and proves that the cost to profit ratio has broke even.
315,000 expense:315,000 profit
315,000-315,000=0